OpenStudy (camilleguidry):

A line passes through the point (–7, 5) and has a slope of one half Which is another point that the line passes through? (–13, 9) (–9, 13) (9, 13) (13, 9)

OpenStudy (welshfella):

Since the slope is 1/2 as the x value increases by 2 the y value increases by 1.

OpenStudy (welshfella):

So for an increase in x of 2 the point (-7,5) goes to (-5,6)

OpenStudy (welshfella):

And if x decreases by 2 the point goes to (-9,4)

OpenStudy (welshfella):

One way to do this would be to subtract the x and y values in (-7,5) from each of the choices If you have the correct choice the the result for the x subtraction will be twice that of the y result.

OpenStudy (mathmale):

"A line passes through the point (–7, 5) and has a slope of one half" translates into y-5 = m(x-[-7]) (this is the "point-slope formula for the equation of a straight line") Thus, y=(1/2)(x+7). Does (9,13) lie on this line? To answer that, let x=9 in the above equation and calculate y. If y = 13, then this point lies on the line. Otherwise, it does not. Continue ....